P1 = 766 torr; T1 = 10.0 °C; P2 = 650.5 torr; T2 = ? °C
1
Expert's answer
2020-05-19T08:46:51-0400

p1/T1= p2/T2

T2= p2 x T1/p1 = 650.5 x 283/ 766 =240.32 K.= -32.68 C